# Joint Discriminative-Generative Modeling via Dual Adversarial Training

### v1

@@ 4.2 Evaluation metrics
We measure both classification and generative modeling performance. For classification, we report clean accuracy and robust accuracy against $L_{2}$ attacks ($\epsilon=0.5$ for CIFAR-10/100 and $\epsilon=3.0$ for ImageNet) computed using AutoAttack (Croce and Hein, 2020). For generative modeling, we evaluate sample diversity and visual fidelity using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016). We focus on conditional generation; details of the generation setup are provided in Section˜A.6.
@@ 4.3 Results
4.3.1 Classification and generative modeling
Comparison with hybrid models. Most existing hybrid models are not explicitly optimized for adversarial robustness, achieving significantly lower robust accuracy than standard AT—for example, JEM (40.5%) and SADA-JEM (31.93%) versus standard AT (75.73%) on CIFAR-10 (Table˜1), and EGC (13.56%) versus standard AT (34.44%/40.28%) on ImageNet (Table˜2). Our DAT approach addresses this limitation by incorporating adversarial training directly into the joint objective, achieving robust accuracy comparable to standard AT. RATIO is the only existing hybrid that explicitly targets robustness, but it does not optimize for generation quality. Our approach achieves comparable robustness to RATIO while substantially improving generative performance. These results demonstrate that our approach bridges the gap between robustness-focused and generation-focused hybrid models, achieving strong performance in both objectives simultaneously. This synergy between robustness and generative capabilities yields practical benefits, as shown in Section˜4.3.2, where our model produces substantially higher-quality counterfactuals through PGD-based generation.
Comparison with generative models. On ImageNet, our DAT approach demonstrates competitive performance relative to dedicated generative models. Our best generative configuration (WRN-50-4 with $T=65$) achieves an FID of 5.39, outperforming BigGAN-deep (6.95) and approaching recent diffusion models such as ADM-G (4.59) and LDM-G (3.60), while requiring significantly less sampling steps. To our knowledge, this represents the first MCMC-based EBM approach to achieve such high-quality generation on complex, high-resolution datasets like ImageNet. Additionally, our model attains relatively strong IS performance, likely due to the PGD-based sampling explicitly optimizing for classifier confidence—a property that aligns with the Inception network-based evaluation metric. We show ImageNet generated samples in Figure˜8.
Trading off generative and discriminative performance.
Our experiments reveal that the number of PGD training steps $T$ in Equation˜8 serves as a mechanism for controlling the balance between discriminative and generative objectives. On CIFAR-10, increasing $T$ from 40 to 50 improves FID from 9.07 to 7.57 at the cost of standard and robust accuracy. A similar trend is observed on CIFAR-100 and ImageNet, where increasing $T$ consistently improves generation quality while reducing classification performance.
@@ 4.3.1 Classification and generative modeling
Effect of model capacity. Our experiments on ImageNet demonstrate the benefits of increased model capacity. Scaling from ResNet-50 (26M parameters) to WRN-50-4 (223M parameters) yields consistent improvements across both discriminative and generative metrics.
Qualitative results. Figures 5, 6, and 7 show generated samples produced by our approach, RATIO, and standard AT. We observe that our method produces visually superior samples with fewer artifacts compared to RATIO and standard AT.
@@
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 hybrid models Residual Flow (Chen et al., 2019) 70.3 3.6 46.4 Glow (Kingma and Dhariwal, 2018) 67.6 3.92 48.9 IGEBM (Du and Mordatch, 2019) 49.1 8.3 37.9 JEM (Grathwohl et al., 2019) 92.9 40.5 8.76 38.4 SADA-JEM (Yang et al., 2023) 95.5 31.93 8.77 9.41 EGC (Guo et al., 2023) 95.9 9.43 3.30 Joint-Diffusion (Deja et al., 2023) 96.4 6.400 RATIO (Augustin et al., 2020) 92.23 76.25 9.61 21.96 Standard AT (Augustin et al., 2020) 92.43 75.73 9.58 28.41 DAT ($T=40$) 91.86 75.66 9.96 9.07 DAT ($T=50$) 90.72 74.65 9.86 7.57 CIFAR-10 conditional generative models SNGAN (Miyato et al., 2018) 8.59 25.5 BigGAN (Brock et al., 2018) 9.22 14.73 StyleGAN2 (Karras et al., 2020a) 9.53 6.96 StyleGAN2 ADA (Karras et al., 2020b) 10.24 3.49 EDM (Karras et al., 2022) 1.79 CIFAR-100 hybrid models Joint-Diffusion (Deja et al., 2023) 77.6 16.8 SADA-JEM (Yang et al., 2023) 75.0 11.63 14.4 EGC (Guo et al., 2023) 77.9 11.50 4.88 RATIO (Augustin et al., 2020) 71.58 47.74 9.28 24.17 Standard AT (Augustin et al., 2020) 72.16 47.78 9.54 23.59 DAT ($T=45$) 65.55 45.97 10.83 10.70 DAT ($T=50$) 62.57 45.97 11.51 9.96
Table 2: Classification and generative modeling results on ImageNet (Standard AT and DAT use $224\times 224$ generation; all other methods use $256\times 256$ generation).
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$ Params Steps Hybrid models Standard AT (Salman et al., 2020) 62.83 34.44 15.97 274.9 26M (ResNet-50) 13 DAT ($T=15$) 57.88 34.84 6.64 339.6 26M (ResNet-50) 13 DAT ($T=30$) 53.46 32.42 5.88 344.0 26M (ResNet-50) 16 Standard AT (Salman et al., 2020) 69.67 40.28 37.25 228.6 223M (WRN-50-4) 11 DAT ($T=30$) 62.09 39.68 6.35 347.6 223M (WRN-50-4) 17 DAT ($T=65$) 54.77 34.14 5.39 341.9 223M (WRN-50-4) 20 EGC (Guo et al., 2023) 78.90 13.56 6.05 231.3 543M 1000 Conditional generative models BigGAN-deep (Brock et al., 2018) 6.95 203.6 340M 1 ADM-G (Dhariwal and Nichol, 2021) 5.44 608M 25 ADM-G (Dhariwal and Nichol, 2021) 4.59 186.7 608M 250 LDM-G (Rombach et al., 2022) 3.60 247.7 400M 250 VAR (Tian et al., 2024) 1.73 350.2 2.0B 10
@@ 4.3.2 Counterfactual generation, OOD detection, and calibration
Counterfactual generation. Figure˜1 compares counterfactual quality across different models while accounting for classifier confidence. Our approach consistently generates counterfactuals with lower FIDs than baseline methods when achieving similar target class confidence. For instance, when the RATIO baseline reaches approximately 0.89 confidence in the target class (at $\epsilon=8$), its corresponding FID is 43.18. Our DAT model achieves a similar confidence level at $\epsilon=4$ with a significantly better FID of 25.53. This demonstrates that, for a comparable level of certainty that the counterfactual represents the target class, our generated samples are substantially more faithful to the true visual characteristics of that class, indicating more plausible counterfactuals. Therefore, our model’s improved generative capability directly translates to higher-quality counterfactual explanations, enhancing model explainability. We provide visualizations of counterfactuals in Section˜A.9.
OOD detection. Our approach generally underperforms RATIO on OOD detection. Ablation studies show this gap persists even when using identical aggressive augmentation for both the generative and discriminative components, indicating it stems from fundamental objective differences rather than the use of milder augmentation for the generative term: RATIO explicitly optimizes for OOD detection while our generative loss prioritizes learning accurate energy functions for generation. The complete details can be found in Section˜A.7
Calibration. Our model’s calibration performance is dataset-dependent, with detailed results provided in Section˜A.8. While the model is well-calibrated on CIFAR-10, outperforming the standard AT and RATIO baselines, it exhibits higher overconfidence on CIFAR-100 and ImageNet. The results suggest that prioritizing generative quality may come at the cost of calibration.
4.3.3 Additional analyses
In Section˜A.5 we conduct additional analyses including component ablation of DAT, OOD dataset effects, and loss weighting mechanisms. Component ablation reveals that both the generative loss and decoupled augmentation contribute to the improved generative quality compared to a standard AT baseline (Section˜A.5.1). OOD dataset analysis demonstrates notable data efficiency of our approach, suggesting strong performance is achievable even with limited auxiliary OOD data (Section˜A.5.2). Loss weighting analysis confirms that re-weighting the two loss terms provides an alternative mechanism for controlling the generative-discriminative trade-off beyond varying PGD steps (Section˜A.5.3).
@@ 5 Conclusion
We addressed the challenge of developing models that excel simultaneously at robust classification and high-fidelity generative modeling. While JEM offers a promising foundation, it suffers from training instability and poor generation quality. Our DAT approach integrates adversarial training principles for both components: replacing unstable SGLD-based JEM learning with an AT-based approach for the generative component, while incorporating standard AT for classification robustness. Experiments across multiple datasets demonstrate that our approach achieves substantially better adversarial robustness over existing hybrid models while maintaining competitive generative capabilities; when optimized for generative performance on ImageNet, it achieves generation quality comparable to dedicated generative models such as GANs and diffusion models. Future work could advance this approach in several directions: improving training efficiency with persistent markov chains, scaling the framework to higher-capacity models to test performance limits; improving secondary tasks like out-of-distribution detection by developing hybrid objectives that combine our generative loss with RATIO’s; and applying the model to broader image synthesis tasks such as those demonstrated by Santurkar et al. (2019).
Appendix A Supplementary Material
A.1 Extended discussion on related work
@@
Learning EBMs with adversarial training Yin et al. [2022] explored an alternative approach to learning EBMs by leveraging the mechanism of Adversarial Training (AT). They established a connection between the objective of binary AT (discriminating real data from adversarially perturbed out-of-distribution data) and the SGLD-based maximum likelihood training commonly used for EBMs. Specifically, they showed that the binary classifier learned via AT implicitly defines an energy function that models the support of the data distribution, assigning low energy to in-distribution regions and high energy to out-of-distribution (OOD) regions. The PGD attack used in AT to generate adversarial samples from OOD data was interpreted as a non-convergent sampler that produces contrastive data, analogous to MCMC sampling in EBM training. Although the resulting energy function can only capture the support rather than recover the exact density, their model achieves competitive image generation performance compared to explicit EBMs. Notably, this AT-based EBM learning approach is more stable than traditional MCMC-based EBM training and demonstrated strong performance in worst-case out-of-distribution detection, similar to methods like RATIO [Augustin et al., 2020].

### v3

@@ 4.2 Evaluation metrics
We measure both classification and generative modeling performance. For classification, we report clean accuracy and robust accuracy against $L_{2}$ attacks ($\epsilon=0.5$ for CIFAR-10/100 and $\epsilon=3.0$ for ImageNet) computed using AutoAttack (Croce and Hein, 2020). For generative modeling, we evaluate sample diversity and visual fidelity using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016). We focus on conditional generation; details of the generation setup are provided in Section˜B.4.
@@ 4.3 Results
We evaluate DAT on CIFAR-10, CIFAR-100, and ImageNet $\times256$ (Tables˜1 and 2). As detailed below, our approach achieves the first successful scaling of EBM-based hybrids to high resolutions while establishing a unique duality between robust classification and high-fidelity generation.
First EBM-based hybrid to scale to high-resolution datasets with adversarial robustness. Prior EBM-based hybrid approaches (JEM, SADA-JEM) are not explicitly optimized for adversarial robustness. On CIFAR-10, these methods achieve significantly lower robust accuracy than standard AT: JEM achieves 40.5% and SADA-JEM achieves 31.93%, compared to 75.73% for standard AT (Table˜1). Our approach addresses this limitation, achieving 75.75% robust accuracy—comparable to standard AT—while improving generative quality over prior EBM hybrids: FID 9.12 versus 38.4 (JEM) and 9.41 (SADA-JEM). Beyond robustness, prior EBM-based hybrids could not scale beyond low resolution or achieve competitive generative performance on ImageNet-level datasets. Our approach is the first to overcome this limitation: on ImageNet $\times256$, our ConvNeXt-Large model achieves FID 3.29 (Table˜2) with classification performance comparable to standard AT (Section˜C.6), demonstrating that EBM-based hybrid models can scale reliably to complex, high-resolution datasets while achieving strong classification performance.
Unique synergy of high-fidelity generation and adversarial robustness. While other scalable hybrids exist, they do not simultaneously achieve both strong adversarial robustness and state-of-the-art generative quality. The diffusion-based EGC achieves 13.56% robust accuracy on ImageNet compared to our 56.40%, while also achieving worse FID (6.05 vs. 3.29) (Table˜2). Similarly, RATIO targets robustness but not generation quality (FID 21.96 vs. our 9.12 on CIFAR-10). Qualitatively, Figures 6, 7, and 8 show that our method produces visually superior samples with fewer artifacts compared to RATIO and standard AT. This unique combination of generative quality and robustness enables critical applications such as robust counterfactual explanations (Section˜4.3.2), where our model produces substantially higher-quality counterfactuals than both non-robust and robustness-only methods.
[…]
4.3.1 Analysis and ablations
Generative-discriminative trade-off. Our experiments reveal that the number of PGD iterations $T$ (Equation˜8) controls the balance between discriminative and generative performance. On CIFAR-10, increasing $T$ from 40 to 50 improves FID from 9.12 to 7.57 at the cost of standard and robust accuracy. A similar trend is observed across all datasets. In Section˜C.7.3, we investigate this inherent tension, revealing that the generative and discriminative objectives encourage different representation alignments. We further show that this trade-off can be explicitly tuned through loss weighting in addition to PGD iterations.
@@ 4.3.1 Analysis and ablations
Effect of model capacity and architecture. Our experiments on ImageNet demonstrate the benefits of increased model capacity and modern architectures. Scaling from ResNet-50 (26M parameters) to WRN-50-4 (223M parameters) yields consistent improvements across both discriminative and generative metrics. Beyond capacity, using state-of-the-art architectures also provides substantial benefits: ConvNeXt-L (198M parameters) substantially outperforms WRN-50-4 (223M parameters) in both accuracy and generation quality despite having fewer parameters, demonstrating the importance of architectural design alongside model scale.
Component contributions. We isolate the contributions of our key components, showing that the generative loss and decoupled augmentation are essential for high-fidelity synthesis (Section˜C.7.1).
OOD data efficiency. We also demonstrate high data efficiency, achieving strong performance even with limited auxiliary OOD data (Section˜C.7.2).
Robustness and stability. Beyond adversarial robustness, DAT maintains strong performance on common corruptions comparable to standard AT (Section˜C.5) and successfully generalizes to $L_{\infty}$ training (Section˜C.6). We also demonstrate high reproducibility with zero training divergence across all runs (Section˜C.8).
[…]
@@
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 hybrid models Residual Flow (Chen et al., 2019) 70.3 3.6 46.4 Glow (Kingma and Dhariwal, 2018) 67.6 3.92 48.9 IGEBM (Du and Mordatch, 2019) 49.1 8.3 37.9 JEM (Grathwohl et al., 2019) 92.9 40.5 8.76 38.4 VERA (Grathwohl et al., 2021) 93.2 8.11 30.5 JEM++ (Yang and Ji, 2021) 94.1 8.11 38.0 JEAT (Zhu et al., 2021) 85.16 8.80 38.24 Robust-JEM (Korst and Asadulaev, 2022) 8.71 41.17 SADA-JEM (Yang et al., 2023) 95.5 31.93 8.77 9.41 WEAT (Mirza et al., 2024) 83.36 8.97 30.74 EGC (Guo et al., 2023) 95.9 9.43 3.30 Joint-Diffusion (Deja et al., 2023) 96.4 6.400 RATIO (Augustin et al., 2020) 92.23 76.25 9.61 21.96 Standard AT (Augustin et al., 2020) 92.43 75.73 9.58 28.41 DAT ($T=40$) 91.92 75.75 9.92 9.12 DAT ($T=50$) 90.72 74.65 9.86 7.57 CIFAR-10 conditional generative models SNGAN (Miyato et al., 2018) 8.59 25.5 BigGAN (Brock et al., 2018) 9.22 14.73 StyleGAN2 (Karras et al., 2020b) 9.53 6.96 StyleGAN2 ADA (Karras et al., 2020a) 10.24 3.49 EDM (Karras et al., 2022) 1.79 CIFAR-100 hybrid models Joint-Diffusion (Deja et al., 2023) 77.6 16.8 SADA-JEM (Yang et al., 2023) 75.0 11.63 14.4 EGC (Guo et al., 2023) 77.9 11.50 4.88 RATIO (Augustin et al., 2020) 71.58 47.74 9.28 24.17 Standard AT (Augustin et al., 2020) 72.16 47.78 9.54 23.59 DAT ($T=45$) 65.76 45.94 10.99 10.73 DAT ($T=50$) 60.12 42.55 11.12 9.53
Table 2: Classification and conditional generative modeling results on ImageNet $\times256$.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$ Params Steps Hybrid models EGC (Guo et al., 2023) 78.90 13.56 6.05 231.3 543M (U-Net) 1000 Standard AT (Salman et al., 2020) 64.91 39.96 15.12 286.2 26M (ResNet-50) 13 DAT ($T=15$) 61.31 39.96 6.87 322.65 26M (ResNet-50) 14 DAT ($T=30$) 55.96 37.14 5.28 319.3 26M (ResNet-50) 14 Standard AT (Salman et al., 2020) 71.25 45.86 37.33 260.2 223M (WRN-50-4) 12 DAT ($T=30$) 64.45 45.84 6.23 341.0 223M (WRN-50-4) 17 DAT ($T=65$) 58.78 40.74 4.94 358.0 223M (WRN-50-4) 19 Standard AT (Singh et al., 2023) 78.25 33.38 44.46 27.32 198M (ConvNext-L-CvSt) 0 DAT ($T=110$) 75.78 56.40 3.29 310.2 198M (ConvNext-L-CvSt) 36 Conditional generative models BigGAN-deep (Brock et al., 2018) 6.95 203.6 340M (ResNet) 1 ADM-G (Dhariwal and Nichol, 2021) 4.59 186.7 608M (U-Net) 250 LDM-4-G (Rombach et al., 2022) 3.60 247.7 400M (U-Net) 250 DiT-XL/2-G (Peebles and Xie, 2023a) 2.27 278.2 675M (Transformers) 250 VAR-d16 (Tian et al., 2024) 3.30 274.4 310M (Transformers) 10 VAR-d30-re (Tian et al., 2024) 1.73 350.2 2.0B (Transformers) 10
@@ 4.3.2 Counterfactual generation, OOD detection, and calibration
Counterfactual generation. Visual counterfactual explanations (VCEs) reveal what minimal semantic changes would flip a classifier’s decision. Standard classifiers cannot generate meaningful VCEs because their input gradients lack semantic structure—gradient-based modifications produce adversarial noise rather than interpretable features. Prior VCE methods were therefore restricted to adversarially robust models (Augustin et al., 2020; Boreiko et al., 2022). DVCE (Augustin et al., 2022) overcomes this limitation by guiding an unconditional diffusion model using gradients from an auxiliary adversarially robust classifier, projecting these onto a cone around the target classifier’s gradient. This enables VCEs for arbitrary classifiers but requires multiple external dependencies: a pretrained diffusion model, a robust classifier, and careful hyperparameter tuning.
In contrast, DAT is intrinsically explainable: it generates VCEs directly by gradient descent on the joint energy $E_{\theta}(x,y)$, requiring no external models. Because our joint objective explicitly learns the energy landscape of the joint distribution $p(x,y)$, the resulting gradients $\nabla_{x}E_{\theta}(x,y)$ naturally point toward semantically valid configurations. Therefore, our model’s improved generative capability directly translates to higher-quality counterfactual explanations.
Figure˜1 compares counterfactual quality across different models while accounting for classifier confidence. Our approach consistently generates counterfactuals with lower FIDs than baseline methods when achieving similar target class confidence. For instance, when the RATIO baseline reaches approximately 0.89 confidence in the target class (at $\epsilon=8$), its corresponding FID is 43.18. Our DAT model achieves a similar confidence level at $\epsilon=4$ with a significantly better FID of 25.53. This demonstrates that, for a comparable level of certainty that the counterfactual represents the target class, our generated samples are substantially more faithful to the true visual characteristics of that class, indicating more plausible counterfactuals. We provide visualizations of counterfactuals in Section˜C.2.
@@
OOD detection. Our approach generally underperforms RATIO on OOD detection. Ablation studies show this gap persists even when using identical aggressive augmentation for both the generative and discriminative components, indicating it stems from fundamental objective differences rather than the use of milder augmentation for the generative term: RATIO explicitly optimizes for OOD detection while our generative loss prioritizes learning accurate energy functions for generation. The complete details can be found in Section˜C.3
Calibration. Our model’s calibration performance is dataset-dependent, with detailed results provided in Section˜C.4. While the model is well-calibrated on CIFAR-10, outperforming the standard AT and RATIO baselines, it exhibits higher overconfidence on CIFAR-100 and ImageNet. The results suggest that prioritizing generative quality may come at the cost of calibration.
@@ 5 Conclusion
We presented Dual Adversarial Training (DAT), demonstrating that hybrid models can achieve state-of-the-art performance without compromise. Our approach advances hybrid modeling in three ways: (1) we present the first EBM-based hybrid to scale to high-resolution complex datasets, achieving FID 3.29 and competitive robust accuracy on ImageNet $\times256$; (2) we uniquely combine state-of-the-art generative quality with adversarial robustness, enabling critical applications such as robust counterfactual explanations; and (3) as a standalone generative model, our approach matches the generative quality of state-of-the-art autoregressive methods and surpasses leading diffusion models while offering unique versatility for diverse synthesis tasks.
Future work could advance this approach in several directions: improving training efficiency with persistent markov chains, scaling the framework to higher-capacity architectures such as ConvNeXt-XLarge and Vision Transformers; improving secondary tasks like out-of-distribution detection by developing hybrid objectives that combine our generative loss with RATIO’s; and applying the model to broader image synthesis tasks such as those demonstrated by Santurkar et al. (2019).
Appendix A Theory & Method
A.1 Extended related work
Connections between adversarial robustness and energy-based models Zhu et al. [2021] reinterpret adversarially trained classifiers as joint energy-based models, showing that adversarial training implicitly flattens the energy landscape around real data by reducing the energy of nearby high-energy adversarial examples. They identify $E_{\theta}(x,y)$ as the key energy term for conditional generation and propose JEAT, which employs energy-based adversarial perturbations and SGLD sampling for likelihood estimation and generation. Mirza et al. [2024] extend this analysis by decomposing the cross-entropy loss as $\mathcal{L}_{\text{CE}}(x,y;\theta)=E_{\theta}(x,y)-E_{\theta}(x)$, revealing that untargeted attacks increase the joint energy $E_{\theta}(x^{*},y)$ (reducing classifier confidence of class $y$) while decreasing the marginal energy $E_{\theta}(x^{*})$. They show that robust overfitting corresponds to divergence between $E_{\theta}(x)$ and $E_{\theta}(x^{*})$, and that state-of-the-art robust models achieve better generalization by smoothing the marginal energy landscape around natural data. Wang et al. [2022b] proposed a unified Contrastive Energy-based Model (CEM) framework that interprets adversarial training as biased maximum likelihood estimation of an energy-based model $p_{\theta}(x,y)=\exp(f_{\theta}(x,y))/Z(\theta)$. Unlike JEM [Grathwohl et al., 2019] which samples negative examples from random noise via Langevin dynamics, CEM shows that PGD-generated adversarial perturbations from real data serve as implicit negative samples, providing more stable training without requiring random noise or OOD data. The framework unifies supervised (P-CEM) and unsupervised (NP-CEM) scenarios, revealing connections between adversarial training, contrastive learning, and energy-based modeling, and enables improved sampling algorithms that achieve state-of-the-art generative performance. While these works provide valuable insights into the energy-based interpretation of adversarial training, training methods that rely on SGLD-based sampling (e.g., JEAT) inherit the instabilities of standard JEM [Grathwohl et al., 2019]. In contrast, our work addresses these stability issues by employing PGD-based contrastive sampling and replacing the unstable EBM loss with a standard BCE loss.
@@ A.1 Extended related work
Learning EBMs with adversarial training Yin et al. [2022] explored an alternative approach to learning EBMs by leveraging the mechanism of Adversarial Training (AT). They established a connection between the objective of binary AT (discriminating real data from adversarially perturbed out-of-distribution data) and the SGLD-based maximum likelihood training commonly used for EBMs. Specifically, they showed that the binary classifier learned via AT implicitly defines an energy function that models the support of the data distribution, assigning low energy to in-distribution regions and high energy to out-of-distribution (OOD) regions. The PGD attack used in AT to generate adversarial samples from OOD data was interpreted as a non-convergent sampler that produces contrastive data, analogous to MCMC sampling in EBM training. Although the resulting energy function can only capture the support rather than recover the exact density, their model achieves competitive image generation performance compared to explicit EBMs. Notably, this AT-based EBM learning approach is more stable than traditional MCMC-based EBM training and demonstrated strong performance in worst-case out-of-distribution detection, similar to methods like RATIO [Augustin et al., 2020]. However, AT-EBM focuses on unconditional generative modeling and employs an explicit $R_{1}$ gradient penalty to stabilize training, which can constrain model expressiveness. Our work incorporates AT-based EBM learning into the JEM framework to perform conditional generative modeling with implicit $R_{1}$ regularization from adversarial training, using ancestral sampling from the conditional distribution $p(x|y)$ rather than the marginal distribution $p(x)$.
Improving joint energy-based models Building on the original JEM framework [Grathwohl et al., 2019], several works have explored techniques to improve training stability and performance. Yang and Ji [2021] (JEM++) introduced multiple training improvements: (1) Proximal SGLD that constrains samples within an $L_{p}$-norm ball of previous samples via gradient clamping for improved stability; (2) YOPO-inspired acceleration (PYLD) that reduces redundant backpropagation by exploiting the coupling between samples and first-layer weights; (3) Informative initialization from a class-conditional Gaussian mixture distribution estimated from training data, which accelerates SGLD convergence, improves stability, and enables batch normalization. Korst and Asadulaev [2022] (Robust-JEM) further enhanced JEM++ by incorporating adversarial training into the discriminative component, empirically observing improved training stability. At inference time, they propose a "combined inference" approach where initial samples from PGD adversarial attacks are refined using SGLD, improving generative performance. However, both JEM++ and Robust-JEM fundamentally still rely on SGLD for sampling and MLE-based objectives for the generative component, inheriting SGLD’s intrinsic instability issues. While these works introduced valuable techniques for improving SGLD stability, they did not fundamentally resolve the instability of SGLD-based training and both remain limited to CIFAR-scale ($\times32$) datasets. Our approach departs fundamentally from this line of work by: (1) providing mathematical analysis and empirical evidence showing that adversarial training offers implicit $R_{1}$ regularization (Section A.3); (2) replacing the MLE-based objective with BCE-based gradients; (3) using deterministic PGD-based sampling instead of stochastic Langevin dynamics; and (4) introducing a two-stage training strategy to address the incompatibility between batch normalization and EBM training. This enables scaling to high-resolution ImageNet synthesis ($\times256$) with state-of-the-art generative quality.